1) Consider two countries, M and N. The value of export of country M is 400 billion US dollar and its GDP is 800 billion US dollar. The value of export of country N is 300 billion US dollar, and its GDP is 900 billion US dollar. Which country has more openness to international trade?
a) Country N
b) Country M
c) It is not possible to know this
d) The two countries are equally open

2) What is the effect of an import tariff imposed by a government on a traded good?
a) Government tariff revenue
b) Change in producer benefits
c) Change in consumer benefits
d) All of the above

3) Suppose that a country imports a product which is also produced by its domestic producers. If tariff is imposed on the import product, what are the effects on domestic producers?
a) It is difficult to know the effect
b) Producer surplus or producer benefits will increase
c) Producer surplus or producer benefits will decrease
d) Domestic producers will get higher price for their product
e) b and d

4) Suppose that company K in the UAE imports 5,000 tonnes of banana from Indonesia for 400 US dollar per tonne. If company K pays an ad-valorem tariff of 10% on the banana imports, how much is the total tariff revenue collected by the UAE government?
a) 500 US dollar
b) 2,000,000 Dirhams
c) 200,000 US dollar
d) 40 US dollar

5) In 2018 UAE imported 200,000 metric tonnes of milk from other countries. The international import price of milk was 1000 US dollars per metric tonne. Domestic consumption of milk in UAE was 250,000 metric tonnes out of which 50,000 metric tonnes was produced in the country. Suppose that the UAE government imposes a 5% tariff on imported milk. The tariff has the following effects: local production increases by 10%; domestic consumption reduces by 10%. After the tariff, how much is the change in domestic consumers’ benefit, the domestic producers’ surplus, the government tax revenue, and deadweight costs, respectively? (use the graph).
a) 2.5 million, 2.65 million, 0.625 million and 0.125 million (in US$)
b) 0.125 million, 0.625 million, 2.5 million and 8.5 million (in US$)
c) -11.875 million, 2.625 million, 8.5 million, and -750,000 (in US$)
d) - 11.875 million, 2.5 million, -0.75 million, and 8.5 million (in US$)
e) 11.875 million, 2.625 million, - 8.5 million, 750,000 (US$)
Answer #1

1.

B

Working note:

Export % of GDP for country M = 400/800 = 50%

Export % of GDP for country N = 300/900 = 33.33%

Since country M has bigger share of international trade, hence country M is more openness to international trade.

2.

D

All the changes as mentioned, will take place.

3.

E

Import tariff will reduce imports and increase the price. At higher price, domestic producers will supply more. So, they will get higher price, also increasing the producer surplus.

4.

C

Working note:

Total tariff revenue = 10% * 400*5000

Total tariff revenue = $200000

Pl. repost other unanswered questions for their proper answers!